# Characterizing Dynamical Stability of Stochastic Gradient Descent in Overparameterized Learning

### v2

@@ 3.6 Stochastic Gradient Descent - the Stable Case
3.7 Generated Matrix Semi-groups
It remains to prove Theorem B (ii). The proof of Theorem B (i) in the previous section was of a quenched nature. One might expect that the best approach to proving Theorem B (ii) is to construct $\omega$-wise center-stable manifolds similar to the ones constructed in section 3.4. While an invariant manifold theory (Pesin theory) has been developed for random dynamical systems (see e.g. Liu and Qian (1995) or Arnold (1998)), these results only provide center-stable manifolds for single points $x\in\mathcal{M}$. For the argument in section 3.4 it was crucial to have a center-stable manifold for an open subset of $\mathcal{M}$. The authors are not aware of any method to construct such a random center-stable manifold.
@@ 3.7 Generated Matrix Semi-groups
We will show this by constructing a Lyapunov function (see e.g. Benaïm and Hurth (2022)) defined on a neighborhood of $x^{*}$ which goes to infinity near $\mathcal{M}$. This is inspired by previous work on the instability of invariant subspaces for stochastic differential equations (cf. Baxendale (1991); Baxendale and Stroock (1988); Bedrossian et al. (2022); Blumenthal et al. (2023); Coti Zelati and Hairer (2021)). While these works rely on conditions of Hörmander-type to establish the existence of a spectral gap in the so-called projective process, such a spectral gap is excluded by the discrete nature of our problem. Instead we use an argument due to Le Page (1982) (cf. also Bougerol and Lacroix (1985) for a survey in English) to find a spectral gap in a different topology. Le Page’s argument needs the matrix semi-group on which the linear co-cycle $\Psi$ is supported to satisfy two algebraic properties, namely being contracting and strongly irreducible (cf. Definition 3.8 below). In this section, we will show that they follow from the regularity of $x^{*}$.
@@
Clearly, $\mathcal{S}(x^{*})$ is a matrix semi-group with unity, generated by $\mathcal{S}_{1}(x^{*})$, i.e. $\mathcal{S}(x^{*})$ contains exactly those matrices, which can be expressed as the product of an arbitrary number of elements in $\mathcal{S}_{1}(x^{*})$, including the empty product, which is defined to be the identity matrix.
@@
A matrix semi-group of invertible matrices $\mathcal{S}\subseteq\operatorname{GL}(N,\mathbb{R})$ is called…
@@ 3.8 Construction of Lyapunov Functions
The goal of this section is, given $\lambda(x^{*})>0$, to construct a Lyapunov function $F^{*}:\mathbb{R}^{N}\setminus\{0\}\to[0,\infty)$ which goes to infinity near 00 and such that
$\mathbb{E}\left[F^{*}\left(\psi_{\omega}^{(1)}(v,w)_{2}\right)\right]\leq\gamma F^{*}(v,w),$
@@
for some $\gamma\in(0,1)$ and every $(v,w)$ in some neighborhood of $(0,0)$. This means that the value of $F^{*}(w)$ must decrease on average along trajectories. Since $F^{*}$ is large near the origin, points get “pushed away” from the set $\{w=0\}$, which will allow us to prove a lack of convergence in the subsequent section.
@@
Following recent works in fluid dynamics (Bedrossian et al. (2022); Blumenthal et al. (2023)), we will study the family of linear operators $(\mathcal{P}_{q}:\mathcal{C}^{0}(S^{N-1})\to\mathcal{C}^{0}(S^{N-1}))_{q\in\mathbb{R}}$ given by
$[\mathcal{P}_{q}f](s)=\mathbb{E}\left[\left\|\Psi^{(1)}_{\omega}s\right\|^{q}f\left(\frac{\Psi^{(1)}_{\omega}s}{\left\|\Psi^{(1)}_{\omega}s\right\|}\right)\right].$
@@
Here $\mathcal{C}^{0}(S^{N-1})$ denotes the Banach space of real-valued continuous functions on $S^{N-1}$, equipped with the supremum-norm. It can be readily seen that the operators $\mathcal{P}_{q}$ are bounded in this norm [^15] In fact $\|\mathcal{P}_{q}\|\leq K_{2}^{q}$ for $q\geq 0$ and $\|\mathcal{P}_{q}\|\leq K_{1}^{-q}$ for $q\leq 0$, with $K_{1}$ and $K_{2}$ defined by (3.28) and (3.29).. The arguments in Bedrossian et al. (2022) and Blumenthal et al. (2023) rely on the fact that $\mathcal{P}_{0}$ admits a spectral gap in their settings. Unfortunately the discrete nature of our setting prohibits such a spectral gap, at least in $\mathcal{C}^{0}(S^{N-1})$. In order to circumvent this obstacle we must consider a different Banach space.
@@
For $\alpha\in(0,1)$, we let $\mathcal{C}^{\alpha}(S^{N-1})$ denote the Banach space of $\alpha$-Hölder continuous functions on the unit sphere $S^{N-1}\subset\mathbb{R}^{N}$, i.e.
@@
Lemma 3.10 (Proposition V.4.1 in Part A of Bougerol and Lacroix (1985))
@@
Lemma 3.10 (ii) says that 1 an dominant eigenvalue of $\mathcal{P}_{0}$, i.e. it is isolated and the spectral value with the largest absolute value. By classical perturbation theory (Kato (1995)) this implies that for $q$ sufficiently close to 0 the largest spectral value of $\mathcal{P}_{q}$ is also an isolated eigenvalue and both the dominant eigenvalue and the corresponding Riesz projections are analytic in $q$. Put more precisely, we get the following corollary (cf. also Theorem V.4.3 in Part A of Bougerol and Lacroix (1985)). Here $\dot{\mathcal{C}}^{\alpha}$ denotes the dual space of $\mathcal{C}^{\alpha}$ and $\mathcal{Q}^{*}\in L(\dot{\mathcal{C}}^{\alpha})$ the dual operator of an operator $\mathcal{Q}\in L(\mathcal{C}^{\alpha})$.
@@
Proof  Note that (3.42) states that $\mathcal{P}_{-p}f^{*}=\gamma f^{*}$. In the following, we will show that $\frac{\mathrm{d}}{\mathrm{d}q}r(q)|_{q=0}=\lambda(x^{*})>0$. This will imply that for sufficiently small $p>0$, we can set $\gamma:=r(-p)<1$ and $f^{*}:=f_{-p}$. Since, for small $p$, the function $f_{-p}$ is close161616in the $\mathcal{C}^{\alpha}$-sense, but thus in particular in the $\mathcal{C}^{0}$-sense to $f_{0}=\mathbf{1}$, it is indeed a positive function and by (3.41), we have $\mathcal{P}_{-p}f^{*}=\gamma f^{*}$.
@@
This establishes that $F^{*}$ is a Lyapunov function for the linearized process induced by $\Psi$. For the remainder of this section, we will show that, on a neighborhood of the origin, $F^{*}$ is also a Lyapunov function for the Markov-process induced by the non-linear cocycle $\psi$. Since $f^{*}$ is continuous, positive and has a compact domain it is both bounded and bounded away from zero, i.e. we can find constants $0<C_{-}\leq C_{+}<\infty$ such that
@@
For every $\varepsilon>0$, there exists a $\delta>0$ s.t.
@@ 3.9 Stochastic Gradient Descent - the Unstable Case
We first argue, why this is sufficient to show $x^{*}\notin\operatorname{supp}(X_{\lim}^{\operatorname{SGD}}).$ Let $x^{*}\in U\subseteq\mathcal{M}$ be the neighborhood given by
@@ Acknowledgements
The authors would like to thank Tim van Erven for his insightful feedback. Furthermore, the authors thank the DFG SPP 2298 for supporting their research. Both authors have been additionally supported by Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689), in the case of D.C. via the Berlin Mathematical School and in the case of M.E. via projects AA1-8, AA1-18 and EF45-5. Furthermore, M.E. thanks the DFG CRC 1114 and the Einstein Foundation for support.
@@ Appendix A Comparison to second moment linear stability
Second moment stability is closely related to so-called moment Lyapunov exponents (see e.g. Arnold (1984)), which have been linked to central limit theorems as well as large deviation theory for the convergence of finite-time Lyapunov exponents (see e.g. Bougerol and Lacroix (1985); Arnold et al. (1986); Arnold and Kliemann (1987)).
@@
To summarize, if $x^{*}\in\mathcal{M}$ is second moment stable in the sense of Wu et al. (2018), it satisfies $\Lambda_{2}(x^{*})\leq 0$ and thus $\lambda(x^{*})\leq 0$. Most of these points should be regular in the sense of Definition 2.2 and even satisfy $\lambda(x^{*})<0$. By Theorem B, these global minima lie in the support of $X_{\lim}^{\operatorname{SGD}}$. On the other hand the one dimensional example demonstrated above suggests that it is possible to have global minima $x^{*}\in\mathcal{M}$ with $\lambda(x^{*})<0$ and $\Lambda_{2}(x^{*})>0$. Given that they are regular, these global minima will be in the support of $X_{\lim}^{\operatorname{SGD}}$, but will not be second moment linearly stable in the sense of Wu et al. (2018). Notably the paper by Wu et al. (2018) contains a sufficient condition for second order linear stability (cf.  Therorem 1 in Wu et al. (2018)). By the argument above, their condition is also a sufficient condition for $\lambda(x^{*})\leq 0$.
In Ma and Ying (2021), a more general notion of “$k$-th order linear stability” is introduced. If $p=k$ is an even natural number this is equivalent to the following generalization of second moment linear stability (cf. Remark 1 in Ma and Ying (2021)).

### v3

@@ 3.6 Stochastic Gradient Descent - the Stable Case
3.7 Generated Matrix semigroups
It remains to prove Theorem B (ii). The proof of Theorem B (i) in the previous section was of a quenched nature. One might expect that the best approach to proving Theorem B (ii) is to construct $\omega$-wise center-stable manifolds similar to the ones constructed in Section 3.4. While an invariant manifold theory (Pesin theory) has been developed for random dynamical systems (see e.g. Liu and Qian, 1995 or Arnold, 1998), these results only provide center-stable manifolds for single points $x\in\mathcal{M}$. For the argument in Section 3.4 it was crucial to have a center-stable manifold for an open subset of $\mathcal{M}$. The authors are not aware of any method to construct such a random center-stable manifold.
@@ 3.7 Generated Matrix semigroups
We will show this by constructing a Lyapunov function (see e.g. Benaïm and Hurth, 2022) defined on a neighborhood of $x^{*}$ which goes to infinity near $\mathcal{M}$. This is inspired by previous work on the instability of invariant subspaces for stochastic differential equations (cf. Baxendale, 1991; Baxendale and Stroock, 1988; Bedrossian et al., 2022; Blumenthal et al., 2023; Coti Zelati and Hairer, 2021). These works rely on conditions of Hörmander type to establish the existence of a spectral gap in the so-called projective process. Such a spectral gap is excluded by the discrete nature of our problem. Instead, we use an argument due to Le Page (1982) (cf. also Bougerol and Lacroix, 1985 for a survey in English) to find a spectral gap in a different topology. Le Page’s argument needs the matrix semigroup on which the linear cocycle $\Psi$ is supported to satisfy two algebraic properties, namely being contracting and strongly irreducible (cf. Definition 3.8 below). In this section, we will show that they follow from the regularity of $x^{*}$.
@@
Clearly, $\mathcal{S}(x^{*})$ is a matrix semigroup with unity, generated by $\mathcal{S}_{1}(x^{*})$, i.e. $\mathcal{S}(x^{*})$ contains exactly those matrices, which can be expressed as the product of an arbitrary number of elements in $\mathcal{S}_{1}(x^{*})$, including the empty product, which is defined to be the identity matrix.
@@
A matrix semigroup of invertible matrices $\mathcal{S}\subseteq\operatorname{GL}(N,\mathbb{R})$ is called,
@@ 3.8 Construction of Lyapunov Functions
The goal of this section is, given a global minimum $x^{*}$ with $\lambda(x^{*})>0$, to construct a Lyapunov function $F^{*}:\mathbb{R}^{N}\setminus\{0\}\to[0,\infty)$ which goes to infinity near 00 and such that
$\mathbb{E}\left[F^{*}\left(\psi_{\omega}^{(1)}(v,w)_{2}\right)\right]\leq\tilde{\gamma}F^{*}(w),$ (3.41)
@@
for some $\tilde{\gamma}\in(0,1)$ and every $(v,w)$ in some neighborhood of $(0,0)$. This means that the value of $F^{*}(w)$ must decrease on average along trajectories. Since $F^{*}$ is large near the origin, points get “pushed away” from the set $\{w=0\}$, which will allow us to prove a lack of convergence in the subsequent section.
To construct $F^{*}$ we employ a strategy that is inspired by recent advances in fluid dynamics (Bedrossian et al., 2022, 2025; Blumenthal et al., 2023). We construct the function $F^{*}$ to be a Lyapunov function for the linearized dynamics, that is,
$\mathbb{E}\left[F^{*}(\Psi_{\omega}^{(1)}w)\right]=\gamma F^{*}(w),~{}\forall\,w\in\mathbb{R}^{N}\setminus\{0\},$ (3.42)
for some $\gamma\in(0,1).$ From this we can conclude (Corollary 3.16 below) that (3.41) holds locally around the origin for any $\tilde{\gamma}\in(\gamma,1)$. Making the ansatz
$F^{*}(w)=\|w\|^{-p}f^{*}\left(\frac{w}{\|w\|}\right),$
with $p>0$ and $f^{*}:S^{N-1}\to[0,\infty)$, equation (3.42) is satisfied if and only if
$\mathbb{E}\left[\left\|\Psi^{(1)}_{\omega}s\right\|^{-p}f^{*}\left(\frac{\Psi^{(1)}_{\omega}s}{\left\|\Psi^{(1)}_{\omega}s\right\|}\right)\right]=\gamma f^{*}(s),~{}\forall\,s\in S^{N-1}.$
@@
We will study the family of linear operators $(\mathcal{P}_{q}:\mathcal{C}^{0}(S^{N-1})\to\mathcal{C}^{0}(S^{N-1}))_{q\in\mathbb{R}}$ given by
$[\mathcal{P}_{q}f](s)=\mathbb{E}\left[\left\|\Psi^{(1)}_{\omega}s\right\|^{q}f\left(\frac{\Psi^{(1)}_{\omega}s}{\left\|\Psi^{(1)}_{\omega}s\right\|}\right)\right].$ (3.43)
@@
Here, $\mathcal{C}^{0}(S^{N-1})$ denotes the Banach space of real-valued continuous functions on $S^{N-1}$, equipped with the supremum norm. It can be readily seen that the operators $\mathcal{P}_{q}$ are bounded in this norm [^16] In fact $\|\mathcal{P}_{q}\|\leq K_{2}^{q}$ for $q\geq 0$ and $\|\mathcal{P}_{q}\|\leq K_{1}^{-q}$ for $q\leq 0$, with $K_{1}$ and $K_{2}$ defined by (3.28) and (3.29).. Furthermore, the operators $\mathcal{P}_{q}$ are positive, that is, they map non-negative functions to non-negative functions. We can reformulate (3.43) as the eigenvalue problem
$\mathcal{P}_{-p}f^{*}=\gamma f^{*},\quad f^{*}\geq 0,~{}0<\gamma<1,~{}p>0.$ (3.44)
Let us pretend for now that the operators $\mathcal{P}_{q}$ are compact and strongly positive [^17] Neither of these properties actually hold., that is, they map any non-negative function that is positive somewhere to a function that is positive everywhere. By the Krein-Rutman Theorem (Krein and Rutman, 1950), for each $q\in\mathbb{R}$, there would be a unique non-negative eigenfunction $f_{q}\geq 0$ whose corresponding eigenvalue, say $r(q)$ is simple, isolated, and principal. In this setting, the principal eigenvalue $r(q)$ is closely related to the $q$-th moment Lyapunov exponent $\Lambda_{q}$ (cf. Definition A.2), by the identity $\Lambda_{q}=\log r(q)$ (see, e.g. Arnold and Kliemann, 1987, Lemma 2). To find a solution to (3.44), we therefore look for a $p>0$ with $\Lambda_{-p}<0$. In settings where the Krein-Rutman Theorem holds for the operators $\mathcal{P}_{q}$, the moment Lyapunov exponent $\Lambda_{q}$ is known to satisfy $\Lambda_{0}=0$ and $\frac{\mathrm{d}}{\mathrm{d}q}\Lambda_{q}|_{q=0}=\lambda$ (Arnold and Kliemann, 1987, Theorem 1). If $\lambda>0$, we can choose $0<p\ll 1$ to have $\Lambda_{-p}<0$ and thus a solution to (3.44).
If the operators $\mathcal{P}_{q}$ are not compact, a similar argument is still possible if one assumes that the operator $\mathcal{P}_{0}$ admits a spectral gap. By definition, the operator $\mathcal{P}_{0}$ is the Markov operator of the linearized dynamics projected onto the unit sphere. In particular, $\|\mathcal{P}_{0}\|=1$ and the constant function $\mathbf{1}\in C^{0}(S^{N-1})$ is an eigenfunction with $\mathcal{P}_{0}\mathbf{1}=\mathbf{1}$. We say that $\mathcal{P}_{0}$ admits a spectral gap if $1$ is a simple eigenvalue and the remaining spectrum is contained in some ball with radius less than 1. This is equivalent to the associated Markov process being uniformly geometrically ergodic (see, e.g. Hairer, 2021). Under the spectral gap assumption one can use tools from perturbation theory (Kato, 1995) to show that the principal eigenvalue $r(q)$ of $\mathcal{P}_{q}$ is analytic in $q$ for a neighborhood of 0. This turns out to be sufficient to find a solution to the eigenvalue problem (3.44), whenever $\lambda>0$ (see Bedrossian et al., 2022, 2025; Blumenthal et al., 2023).
@@
Unfortunately, the discrete nature of our setting prohibits uniform geometric ergodicity and thus a spectral gap for $\mathcal{P}_{0}$, at least in $\mathcal{C}^{0}(S^{N-1})$. In order to circumvent this obstacle, we must consider a different Banach space. For $\alpha\in(0,1)$, we let $\mathcal{C}^{\alpha}(S^{N-1})$ denote the Banach space of $\alpha$-Hölder continuous functions on the unit sphere $S^{N-1}\subset\mathbb{R}^{N}$, i.e.
@@
Lemma 3.10 (Proposition V.4.1 in Part A of Bougerol and Lacroix, 1985)
@@
Lemma 3.10 (ii) says that 1 is an dominant eigenvalue of $\mathcal{P}_{0}$, i.e. it is isolated and the spectral value with the largest absolute value. By classical perturbation theory (Kato, 1995) this implies that for $q$ sufficiently close to 0 the largest spectral value of $\mathcal{P}_{q}$ is also an isolated eigenvalue and both the dominant eigenvalue and the corresponding Riesz projections are analytic in $q$. Put more precisely, we get the following corollary (cf. also Theorem V.4.3 in Part A of Bougerol and Lacroix, 1985). Here $\dot{\mathcal{C}}^{\alpha}$ denotes the dual space of $\mathcal{C}^{\alpha}$ and $\mathcal{Q}^{*}\in L(\dot{\mathcal{C}}^{\alpha})$ the dual operator of an operator $\mathcal{Q}\in L(\mathcal{C}^{\alpha})$.
@@
Proof  Note that (3.46) states that $\mathcal{P}_{-p}f^{*}=\gamma f^{*}$. In the following, we will show that $\frac{\mathrm{d}}{\mathrm{d}q}r(q)|_{q=0}=\lambda(x^{*})>0$. This will imply that for sufficiently small $p>0$, we can set $\gamma:=r(-p)<1$ and $f^{*}:=f_{-p}$. Since, for small $p$, the function $f_{-p}$ is close181818in the $\mathcal{C}^{\alpha}$-sense, but thus in particular in the $\mathcal{C}^{0}$-sense to $f_{0}=\mathbf{1}$, it is indeed a positive function and by (3.45), we have $\mathcal{P}_{-p}f^{*}=\gamma f^{*}$.
@@
This establishes that $F^{*}$ is a Lyapunov function for the linearized process induced by $\Psi$. For the remainder of this section, we will show that, in a neighborhood of the origin, $F^{*}$ is also a Lyapunov function for the Markov process induced by the nonlinear cocycle $\psi$. Since $f^{*}$ is continuous, positive, and has a compact domain, it is both bounded and bounded away from zero, i.e. we can find constants $0<C_{-}\leq C_{+}<\infty$ such that
@@
For every $\varepsilon>0$, there exists a $\delta>0$ such that
@@ 3.9 Stochastic Gradient Descent - the Unstable Case
We first argue why this is sufficient to show $x^{*}\notin\operatorname{supp}(X_{\lim}^{\operatorname{SGD}}).$ Let $x^{*}\in U\subseteq\mathcal{M}$ be the neighborhood given by
@@ Acknowledgements
The authors would like to thank Tim van Erven for his insightful feedback. Furthermore, the authors thank the DFG SPP 2298 for supporting their research. Both authors have been additionally supported by Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689), in the case of D.C. via the Berlin Mathematical School and in the case of M.E. via projects AA1-8, AA1-18 and EF45-5. Furthermore, M.E. thanks the DFG CRC 1114, the Einstein Foundation and the Dutch Research Council NWO (VI.Vidi.233.133) for support.
@@ Appendix A Comparison to second moment linear stability
Second moment stability is closely related to so-called moment Lyapunov exponents (see e.g. Arnold, 1984), which have been linked to central limit theorems as well as large deviation theory for the convergence of finite-time Lyapunov exponents (see e.g. Bougerol and Lacroix, 1985; Arnold et al., 1986; Arnold and Kliemann, 1987).
@@
To summarize, if $x^{*}\in\mathcal{M}$ is second moment stable in the sense of Wu et al. (2018), it satisfies $\Lambda_{2}(x^{*})\leq 0$ and thus $\lambda(x^{*})\leq 0$. Most of these points should be regular in the sense of Definition 2.2 and even satisfy $\lambda(x^{*})<0$. By Theorem B, these global minima lie in the support of $X_{\lim}^{\operatorname{SGD}}$. On the other hand the one dimensional example demonstrated above suggests that it is possible to have global minima $x^{*}\in\mathcal{M}$ with $\lambda(x^{*})<0$ and $\Lambda_{2}(x^{*})>0$. Given that they are regular, these global minima will be in the support of $X_{\lim}^{\operatorname{SGD}}$, but will not be second moment linearly stable in the sense of Wu et al. (2018). Notably the paper by Wu et al. (2018) contains a sufficient condition for second order linear stability (cf.  Therorem 1 in Wu et al., 2018). By the argument above, their condition is also a sufficient condition for $\lambda(x^{*})\leq 0$.
In Ma and Ying (2021), a more general notion of “$k$-th order linear stability” is introduced. If $p=k$ is an even natural number this is equivalent to the following generalization of second moment linear stability (cf. Remark 1 in Ma and Ying, 2021).